Delta G for acetic acid dissociation is 27.1 kj/mol. calculate the Ka for acetic acid at 25 degrees C.
Answer #1

Sol.

As Standard change in gibbs free energy = deltaG°

= 27.1 KJ/mol = 27100 J/mol  

And , deltaG° = - R × T × ln(Ka)

where R = Gas constant = 8.314 J/ K mol

T = Temperature = 25°C = 298 K  

Ka = Acetic acid dissociation constant

So , 27100 = - 8.314 × 298 × ln(Ka)  

ln(Ka) = - 10.9381

Therefore , Ka = e-10.9381 =   1.77 × 10-5